OpenStudy (compassionate):

What if we have fractions, @rational For example: f(x) = 12/4x+2. Find f(-1) So, we're finding the inverse, too, yes? I would basically isolate x and do the same thing, correct?

OpenStudy (rational):

Exactly ! but use proper parenthesis so that the expression is not ambiguous

OpenStudy (compassionate):

y = 12/4x+2 y * 12 = 4x + 2 y * 12 - 2 = 4x y * 12 -2/4 = x y * 10/4 = x Are we off to a good start?

OpenStudy (rational):

12/4x+2 do you mean \[\large \dfrac{12}{4x}+2\] or \[\large \dfrac{12}{4x+2}\] or \[\large \dfrac{12}{4}x+2\] or what ?

OpenStudy (compassionate):

b

OpenStudy (rational):

i thought so, you should use parenthesis to group the bottom stuff : f(x) = 12/(4x+2)

OpenStudy (rational):

y = 12/(4x+2) cross multiply : 4x+2 = 12/y subtract 2 both sides : 4x = (12/y) - 2 divide 4 both sides : x = (3/y) - 1/2

hartnn (hartnn):

you need f(-1) or \(f^{-1}\) ?

hartnn (hartnn):

f(-1) is just 1 plugged in place of x

OpenStudy (compassionate):

Yes, it's f(-1) not f^-1, but what does f^-1 mean?

OpenStudy (anonymous):

@Compassionate if it is confusing, solve \[\large \dfrac{12}{4x+2}=10\] and see the method

OpenStudy (anonymous):

then solve \[\large \dfrac{12}{4x+2}=\pi\]

OpenStudy (compassionate):

Okay, so: x = (3/y) - 1/2 I understand how we got this, now we just need to plug in -1, right?

OpenStudy (anonymous):

and finally solve \[\large \dfrac{12}{4x+2}=y\] or if you prefer solve \[\large \dfrac{12}{4y+2}=x\]

hartnn (hartnn):

\(f^{-1}\) is the inverse function of f f(-1) is the value of f(x) when x =-1

OpenStudy (anonymous):

\[f^{-1}(x)\] has nothing to do with the number \(-1\) it means the inverse function

OpenStudy (compassionate):

Okay, Satellite, If I do the last one, we have. 12/(4x+2) = y 12/y = 4x+2 10/y = 4x (10/y)/4 = x

OpenStudy (anonymous):

on the other hand, i see that you wrote find \(f(-1)\) which is a whole different story than find \(f^{-1}(x)\)

OpenStudy (anonymous):

which is it?

OpenStudy (compassionate):

It's f(-1), not hte inverse.

hartnn (hartnn):

then don't find the inverse ! :P just plug in x=-1 in f(x)

hartnn (hartnn):

f(-1) = 12/ (-4+2) = ...

OpenStudy (anonymous):

\[\large f(x)=\frac{12}{4x+2}\\ f(-1)=\frac{12}{4\times (-1)+2}\]

hartnn (hartnn):

\(\large f(x)=\frac{12}{4x+2}\\ f(-1)=\frac{12}{4\times (-1)+2} \\ f(\heartsuit )=\frac{12}{4\times (\heartsuit )+2}\)

OpenStudy (anonymous):

hey!

OpenStudy (anonymous):

i patented \(f(\clubsuit)\)

hartnn (hartnn):

lol, thats why i was careful not to use it :P

OpenStudy (compassionate):

Okay, so... f(x) = 12/(4x+2). Find f(-1) y = 12/(4[-1]+ 2) y = 12/(-4 + 2) y = 12/-2 y = -6

hartnn (hartnn):

\(\huge \checkmark\)

OpenStudy (compassionate):

It's the idea of an inverse that throws me off. Could you explain a bit more in detail on this?

hartnn (hartnn):

as sat said, inverse is altogether a different story! f(-1) and \(f^{-1}\)

hartnn (hartnn):

so, you want details about inverse or f(-1) ?

OpenStudy (compassionate):

f^-1

OpenStudy (compassionate):

\[f(x) = 7x - 13. Find f^{-1}(x).\] For example

hartnn (hartnn):

that is inverse function and has nothing to do with this question

OpenStudy (compassionate):

I am aware, but I am asking for you to elaborate a little bit on an inverse and what it is. Would you prefer me to close and open a new question?

hartnn (hartnn):

a function f and g are inverses of each other, iff \(f(g(x)) =x , \\ and \\ g(f(x))=x\) to get inverse function of y=f(x) , we replace x by y and y by x and then isolate x this is essentially to make f(g(x)) =x

OpenStudy (compassionate):

f(x) = 7x - 13. Find f^-1(x) So I could say. y = 7x - 13 y + 13 = 7x (y + 13)/7 = x And then reverse (x + 13)/7 = y, Right?

hartnn (hartnn):

absolutely! :)

OpenStudy (compassionate):

I'm really getting the hang of this (:

hartnn (hartnn):

cool :) find the inverse of 1/x :)

OpenStudy (compassionate):

*dies*

OpenStudy (compassionate):

1/x isn't an equation. How does one find the inverse.

hartnn (hartnn):

ok then f(x) =1/x

OpenStudy (compassionate):

f(x) = 7x - 13. Find f(x) = 1/x Uh. f(x) = 7(1/x) - 13? cx

hartnn (hartnn):

no...i gave you another function.... forget that f(x) was 7x-13

hartnn (hartnn):

ok, g(x) =1/x find \(g^{-1}(x)\)

OpenStudy (compassionate):

Uh. y = 1/x y * 1 = x? x * 1 = y

hartnn (hartnn):

so whats xx ?

hartnn (hartnn):

x***

OpenStudy (compassionate):

g(x) =1/x xx? No no. I solve for x. I replaced g(x) with y. y = 1/x. I undid the division by multiplying So y * 1 = x Then I "inversed" them by swapping y and x.

hartnn (hartnn):

ok, so whats your final answer for the inverse function ?

OpenStudy (compassionate):

It's x * 1 = x

hartnn (hartnn):

inverse function is a function \(g^{-1}(x) =... ?\) (in terms of x)

OpenStudy (compassionate):

My brain has reached maximum capacity.

hartnn (hartnn):

x=1/y so y=1/x so \(g^{-1}(x) = 1/x\) an example of the function whose inverse is itself

OpenStudy (compassionate):

This is too much. My brain wasn't built for this.